(1). Find the compound interest on 10,000 for 12 months at rate of 1% peq month. (2) if x varies directly as y and x=32 at y=2 find the relationshi betwee x and y . Also x when y=4 and y when x=64 (3) x-3 is directly proportional to the squere of y and x=5 when y=2, find x when y=6 (my gmail mohdbukar4@gmail.com)
Answer by Boreal(15235)

If the interest rate is 1%/month then 10000*(1+0.01)^12=11,268.05
====================
x=ky
32=k*2
k=16
x=16y
-------------
x=16*4=64 so when y is 4, x is 64
==========
(x-3)=ky^2
(5-3)=k*4
2=4k
k=0.5
==========
when y=6
(x-3)=0.5*36=18
x=21 when y=6
(5,2) and (21, 6); take 3 away from x and 2 is proportional to 4 as 18 is proportional to 36.